OpenStudy (anonymous):

How do i find the axis of symmetry of the equation y=a(X-5.68)^2+5.1

OpenStudy (anonymous):

This equation is in Vertex form: y=a(x-h)^2 + k Do you know how to find the vertex from this equation? (vertex is the maximum of the parabola)

OpenStudy (anonymous):

It is important to learn that the axis of symmetry lies on the vertex. The axis of symmetry is represented by the equation x=? Where x is the x coordinate of the vertex.

OpenStudy (anonymous):

From the given equation, you can simply read this off as 5.68. Be sure to include the correct sign though. If the equation was (X+5.68), then the equation for the axis of symmetry would be x= -5.86

OpenStudy (anonymous):

Okay and how do i put that equation into y=ax^2+bx+C?

OpenStudy (anonymous):

If i substitute x for 1

OpenStudy (anonymous):

First you deal with the exponent which is the ^2. So now, using perfect squares expand.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Do you know what 'a' is?

OpenStudy (anonymous):

No. Thats why im confused. What is "a" in the equation

OpenStudy (anonymous):

Hang on, I think I know how to find it - im working on it. 'a' controls how skinny or fat the parabola/curve is.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

One of the ppints that the parabola pass through is (9,2) so i would have 2=a(9-5.68)^2+5.1?

OpenStudy (anonymous):

To find a, you need a point (x,y) that lies on the equation. Has this been given to you? Is there any other information that you know?

OpenStudy (anonymous):

Yep, you got it! :)

OpenStudy (anonymous):

& i got -0.28?

OpenStudy (anonymous):

I got: 2=a(9-5.68)62 +5.1 -3.1=a(3.32)^2 -3.1 = a11.0224 therefore, a = -3.1/11.0224 a=-0.281

OpenStudy (anonymous):

Well done :)

OpenStudy (anonymous):

okay. Me too so now how i do put that into the y=ax^2+bx+c thing

OpenStudy (anonymous):

My equation is y=-0.28(x-5.68)^2+5.1

OpenStudy (anonymous):

First step - expand (x-5.68)2 using perfect squares do you know how to do this?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok hang on i'll write it out

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

im about to upload my working - wont take a sec

OpenStudy (anonymous):

Do you have an answer?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Am i solving for x?

OpenStudy (anonymous):

Just to let you now that the vertex, which is your axis of symmetry can be easily found by nowing that the h = x and k = y in the equation \(a(x-h)^2 + k \) You equation \( y=a(X-5.68)^2+5.1 \) x = -5.68 y = 5.1 The vertex and your axis of symmetry = (-5.68, 5.1) That is your point and draw a line straight up and that will be your axis of symmetry

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

No you are not solving for x - you are expanding into general form - isn't that what you asked?

OpenStudy (anonymous):

Yes, i got the -0.28x^2+3.1808x-3.933

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thats right?

OpenStudy (anonymous):

Yes - i got the same as you :)

OpenStudy (anonymous):

okay thanks alot!

OpenStudy (anonymous):

One more question i have to find the discriminant ? How do i do that

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

Okay, so do you know what the equation for the discriminant is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

it is\[\sqrt{b ^{2}-4ac}\]

OpenStudy (anonymous):

Okay hold on

OpenStudy (anonymous):

find a, b and c from the equation you just found (general form) and calculate

OpenStudy (anonymous):

Okay so its the sqrt of 3.1808^2-4(-0.28)(3.933)?

OpenStudy (anonymous):

I got that the discriminant equals 2.39... Therefore since \[2.39>0\] then the quadratic has two distinct real roots In other words, it cuts the x-axis twice

OpenStudy (anonymous):

yes - everything is under the square root

OpenStudy (anonymous):

i got 3.81

OpenStudy (anonymous):

Yes you are right - just did it again sorry

OpenStudy (anonymous):

So do i find the x-intercepts from that?

OpenStudy (anonymous):

okay so its 3.81>0?

OpenStudy (anonymous):

yes. 2.39>0 then the quadratic has two distinct real roots In other words, it cuts the x-axis twice. Does the question require you to find the x intercepts?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The x & y-intercepts

OpenStudy (anonymous):

The discriminant does not tell you what the x - intercepts are, it simply indicates if there are two roots (solutions), d\[\Delta \] which is the discriminant , if \[\Delta>0\]two solutions if \[\Delta=0\]one solution if \[\Delta<0\]no real roots (complex roots)

OpenStudy (anonymous):

Anyway, to find the x intercept just let y=0

OpenStudy (anonymous):

To find y intercepts, let x=0

OpenStudy (anonymous):

And for the y?

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

So its 0=-0.28(x-5.68)^2+5.1? & y=-0.28(0-5.68)^2+5.1?

OpenStudy (anonymous):

yep - just solve now

OpenStudy (anonymous):

i got x intercepts are=9.95,1.41 & y intercepts are=-3.93

OpenStudy (anonymous):

An x intercept should have co-ordinates (x,0) and y-in = (0,y) try again

OpenStudy (anonymous):

You should get two x intercepts, as indicated by the discriminant

OpenStudy (anonymous):

so x ints are (1.41,0) and (9.95,0) y int is at (0,-3.93) you are correct

OpenStudy (anonymous):

Yes thats what i got

OpenStudy (anonymous):

:)